OpenStudy (anonymous):

how can you know the sum and difference of 2 cubes are being factored ?

OpenStudy (anonymous):

what do you mean, "how can you know"? If you have a sum of 2 cubes, or a difference of two cubes, it can always be factored.

OpenStudy (anonymous):

i cant understand it .. this the the question ... know how the sum and difference of 2 cubes are being factored ..

OpenStudy (anonymous):

do you have a problem you are trying to work out? the two formulas you need to know are: \[x^3-y^3 = (x-y)(x^2+xy+y^2)\] and \[x^3+y^3 = (x+y)(x^2-xy+y^2)\]

OpenStudy (anonymous):

none .. just a question :/ math is hard for me :'(

OpenStudy (anonymous):

alright, well here is an example: \[x^3-125\] Lets say you want to factor that. First you would need to notice that: \[125 = 5^3\] so we can re-write the question like this: \[x^3-125 \Rightarrow x^3-5^3\] Then, you should be thinking, "this is a difference of 2 cubes, i need to use that formula!" So you look through your note and find the formula: \[x^3-y^3 = (x-y)(x^2+xy+y^2)\] but how do you use it? put the left sides of the equation close to each other: \[x^3-y^3\] \[x^3-5^3\] it looks like the 'x' stayed the same, but the y changed to a 5. So in the long expression: \[(x-y)(x^2+xy+y^2)\] change all the y's to 5's, and simplify if you can: \[(x-5)(x^2+x(5)+5^2) = (x-5)(x^2+5x+25)\] and then you are done :) so the answer would be: \[x^3-5^3 = (x-5)(x^2+5x+25)\] That is how you factor a difference of 2 cubes.

OpenStudy (anonymous):

thanks to you .... thanks for your help im so very happy :) thank you and god bless more answer to come :) peace ....